OpenStudy (anonymous):

How do you find the derivatives? int_{0}^{4t^2}sqrt(1-3x^3)dt int_{3}^{lnx}tan(t) dt

OpenStudy (anonymous):

The fundamental theorem of calculus \[ F(x) = \int_a^x f(t)dt \implies F'(x) = f(x) \]

OpenStudy (anonymous):

for the first one, do you mean to find F'(x) when \(\large F(x)=\int_{0}^{4x^2}\sqrt{1-3t^3}dt \) ???

OpenStudy (anonymous):

Yes. That's what I mean. I don't really know how to use that theorem.

OpenStudy (anonymous):

\[ \large F(x)=\int_{0}^{4x^2}\sqrt{1-3t^3}dt \ \]So we can't quite use the fundamental theorem yet, because of that nasty \(4x^2\). Let's suppose that: \(y=4x^2\) and \(F(x) = G(y) \). Using the fundamental theorem, we get: \[ \large G(y)=\int_{0}^{y}\sqrt{1-3t^3}dt \implies G'(y) = \sqrt{1-3y^3} \]Now, consider that \(F(x) = G(y) = G(y(x))\). If we use the chain rule, we get: \[ \large F'(x) = G'(y(x)) \cdot y'(x) \]

OpenStudy (anonymous):

We know \[ \large G'(y(x)) = \sqrt{1-3[y(x)]^3} =\sqrt{1-3[4x^2]^3} = \sqrt{1-192x^6} \\ \large y(x) = 4x^2 \\ \large y'(x) = 8x \\ \large F'(x) = \sqrt{1-192x^6} \cdot 8x \]

OpenStudy (anonymous):

For problem 2, we get: \[ G'(y) = \tan(y) \\ y'(x) = \frac{1}{x} \\ F'(x) = \frac{\tan(\ln(x))}{x} \]